### v1

# SVTRv2: CTC Beats Encoder-Decoder Models in Scene Text Recognition

Abstract
Connectionist temporal classification (CTC)-based scene text recognition (STR) methods, e.g., SVTR, are widely employed in OCR applications, mainly due to their simple architecture, which only contains a visual model and a CTC-aligned linear classifier, and therefore fast inference. However, they generally have worse accuracy than encoder-decoder-based methods (EDTRs), particularly in challenging scenarios. In this paper, we propose SVTRv2, a CTC model that beats leading EDTRs in both accuracy and inference speed. SVTRv2 introduces novel upgrades to handle text irregularity and utilize linguistic context, which endows it with the capability to deal with challenging and diverse text instances. First, a multi-size resizing (MSR) strategy is proposed to adaptively resize the text and maintain its readability. Meanwhile, we introduce a feature rearrangement module (FRM) to ensure that visual features accommodate the alignment requirement of CTC well, thus alleviating the alignment puzzle. Second, we propose a semantic guidance module (SGM). It integrates linguistic context into the visual model, allowing it to leverage language information for improved accuracy. Moreover, SGM can be omitted at the inference stage and would not increase the inference cost. We evaluate SVTRv2 in both standard and recent challenging benchmarks, where SVTRv2 is fairly compared with 24 mainstream STR models across multiple scenarios, including different types of text irregularity, languages, and long text. The results indicate that SVTRv2 surpasses all the EDTRs across the scenarios in terms of accuracy and speed. Code is available at https://github.com/Topdu/OpenOCR.
1 Introduction
As a task of extracting text from natural images, scene text recognition (STR) has garnered considerable interest over decades. Unlike text from scanned documents, scene text often exists within complex natural scenarios, posing challenges such as background noise, text distortions, irregular layouts, artistic fonts [7], etc. To tackle these challenges, a variety of STR methods have been developed and they can be roughly divided into two categories, i.e., connectionist temporal classification (CTC)-based methods and encoder-decoder-based methods (EDTRs).
Typically, CTC-based methods [39, 11, 23, 28] employ a single visual model to extract image features and then apply a CTC-aligned linear classifier [16] to predict recognition results. This straightforward architecture provides advantages such as fast inference, which makes them especially popular in OCR applications. However, these models struggle to handle text irregularity, i.e., text distortions, varying layouts, etc. As a consequence, advanced attention-based decoders are introduced as alternatives to the CTC classifier, leading to a series of EDTRs [40, 38, 32, 29, 47, 55, 36, 59, 57, 15, 48, 4, 34, 46, 12, 58, 8, 62, 54, 49, 52, 9, 51, 17, 18, 63]. These attention-based decoders show appealing performance in integrating multi-modal cues, including visual [48, 58, 12, 52], linguistic [38, 55, 15, 36], and positional [57, 62, 8] ones, which are largely missed in current CTC models. This integration enables EDTRs to perform more effectively in complex scenarios. As depicted in the top of Fig. 1, compared to SVTR [11], a leading CTC model that is adopted by famous commercial OCR engines [28], EDTRs achieve superior results in English and Chinese benchmarks [25, 6], covering challenging scenarios such as curved, multi-oriented, occluded, and artistic text.
Nevertheless, EDTRs are commonly built upon complex architectures, thus sacrificing the inference speed (FPS), as shown in the top of Fig. 1. In addition, besides slower inference speed, EDTRs do not handle long text well. Even LISTER [8], an EDTR dedicated to long text recognition, performs worse than SVTR [11]. Since fast response and recognizing long text are both important for many applications, the OCR community has to face the dilemma that no model excels in accuracy, speed and versatility. When selecting either CTC-based models or EDTRs, users have to accept that the model is inferior in some aspects.
The inferior accuracy of CTC models can be attributed to two primary factors. First, these models struggle with irregular text, as CTC alignment presumes that the text appears in a near canonical left-to-right order [7, 2], which is not always true, particularly in complex scenarios. Second, CTC models seldom encode linguistic information, which is typically accomplished by the decoder of EDTRs. While recent advancements deal with the two issues by employing text rectification [40, 32, 61], developing 2D CTC [44], utilizing masked image modeling [48, 58], etc., the accuracy gap between CTC and EDTRs remains significant, indicating that novel solutions still need to be investigated.
In this paper, we aim to build more powerful CTC models by better handling text irregularity and integrating linguistic context. For the former, we address this challenge by first extracting discriminative features and then better aligning them. First, existing methods uniformly resize text images with various shapes to a fixed size before feeding into the visual model. We question the rationality of this resizing, which easily causes severe distortion of the text, making it unreadable to humans, as shown in the bottom-left of Fig. 1. To this end, we propose a multi-size resizing (MSR) strategy to adaptively resize text images according to their aspect ratios, thus minimizing text distortion and ensuring the discrimination of the extracted visual features. Second, irregular text may be rotated significantly, and the character arrangement does not align with the reading order of the text, causing the puzzle for CTC alignment, as shown in the bottom-centre example in Fig. 1. To solve this, we introduce a feature rearrangement module (FRM) which rearranges visual features with a horizontal rearrangement, and then identifying and prioritizing relevant vertical features. FRM maps 2D visual features into a sequence aligned with the text’ reading order, thus effectively alleviating the alignment puzzle. Consequently, CTC models integrating MSR and FRM can recognize irregular text effectively, without using rectification modules or attention-based decoders.
As for the latter, the mistakenly recognized example shown in the bottom-right of Fig. 1 clearly highlights the necessity of integrating linguistic information. Since CTC models directly classify visual features, we have to endow the visual model with linguistic context modeling capability, which is less discussed previously. To this end, inspired by guided training of CTC (GTC) [23, 28] and string matching-based recognition [13], we propose a semantic guidance module (SGM), which devises a new scheme that solely leverages surrounding string context to recognize target characters during training. This approach effectively guides the visual model to learn to perceive the linguistic context without rely on decoders. During inference, SGM can be omitted and would not increase the inference cost.
With these contributions, we develop SVTRv2, a novel CTC-based method whose recognition ability has been largely enhanced, while still maintaining a simple architecture and fast inference. To thoroughly validate SVTRv2, we conducted extensive ablation and comparative experiments on benchmarks including standard regular and irregular text [2], occluded scene text [48], the recent Union14M-L benchmarks [25], long text [13], and Chinese [6] text. The results demonstrate that SVTRv2 consistently outperforms all the compared EDTRs across the evaluated scenarios in terms of accuracy and speed, highlighting its effectiveness and broad applicability.
In addition, recent advances [4, 25, 37] revealed the importance of large-scale real-world datasets in improving STR model performance. However, many STR methods primarily derived from synthetic data [19, 24], which fail to fully represent real-world complexities and lead to performance limitations, particularly on challenging scenarios. Notably, we observe that the existing real-word datasets [4, 25, 37] suffer from data leakage, and the results reported in [25] should be updated. As a result, we introduce U14M-Filter, a rigorously filtered version of the real-world training dataset Union14M-L [25]. We systematically reproduced and retrained 24 mainstream STR methods from scratch based on U14M-Filter. These methods are also thoroughly evaluated across various STR benchmarks. Their accuracy, model size, and inference time constitute a comprehensive and reliable new benchmark for future reference.
2 Related Work
Irregular text recognition [1, 35, 25] has posed a significant challenge in STR due to the diverse variation of text instances, where CTC-based methods [39, 11, 23, 28] are often less effective. To address this, some methods [40, 36, 59, 62, 54, 11, 61] incorporate rectification modules [40, 32, 61] that aim to transform irregular text into more regular format. Alternatively, more methods utilize attention-based decoders [47, 29, 38], which employ the attention mechanism to dynamically localize characters regardless of text layout, and thus less affected. However, these methods generally have tailored training hyper-parameters. For instance, the rectification modules [40, 32, 61] typically specify a fixed output image size (e.g. 32128), which is not always a suitable choice. While attention-based decoders [47, 29, 38] generally set the maximum recognition length to 25 characters, thus longer text cannot be correctly recognized, as shown in Fig. 5.
Linguistic Context Modeling. There are several ways of modeling linguistic context. One major branch is auto-regressive (AR)-based STR methods [40, 47, 38, 29, 57, 25, 51, 62, 52, 54, 63, 14], which utilize previously decoded characters to model contextual cues. However, their inference speed is slow due to the character-by-character decoding nature. Some other methods [55, 15, 34, 4] integrate external language models to model the linguistic context and correct the recognition results. While effective, the linguistic context is purely text-based, making it challenging to adapt them to CTC models. There also are some studies [36, 48, 58] to modeling linguistic context with visual information only by using masked image modeling-based pretraining [21, 3]. However, they still depend on attention-based decoders to unleash the linguistic information, not integrating linguistic cues into the visual model, thus limiting their effectiveness in enhancing CTC models.
3 Methods
Fig. 2 illustrates the overview of SVTRv2. A text image is first resized by MSR to the closest aspect ratio, forming the input X . X then experiences three consecutive feature extraction stages, yielding visual features F . During training, F is fed into both SGM and FRM. SGM guides SVTRv2 to model linguistic context, while FRM rearranges F into the character feature sequence , which is synchronized with the text reading order and aligns with the label sequence well. During inference, the SGM is discarded for efficiency.
3.1 Multi-Scale Resizing
Previous works typically resize irregular text images to a fixed size, such as , which, however, may cause undesired text distortion and severely affect the quality of extracted visual features. To address this issue, we propose a simple yet effective multi-size resizing (MSR) strategy that resizes text shapes based on the aspect ratio (R=). Specifically, we define four specific sizes: [64, 64], [48, 96], [40, 112], and [32, 32], respectively corresponding to aspect ratio: R1.5 (R1), 1.5 R2.5 (R2), 2.5R3.5 (R3), and R3.5 (R4). Therefore, MSR allows text instances adaptively resized under the principles of roughly maintaining their aspect ratios, such that significant text distortion caused by resizing is almost eliminated. As a result, the quality of extracted visual features is guaranteed.
3.2 Visual Feature Extraction
Motivated by SVTR [11], the network architecture of SVTRv2 comprises three stages, with containing mixing blocks, as illustrated in Fig. 2. To extract discriminative visual features, we devise two types of mixing blocks: local and global. Local mixing is implemented through two consecutive grouped convolutions, which are expected to capture local character features, such as edges, textures, and strokes. Meanwhile, global mixing is realized by the multi-head self-attention (MHSA) mechanism [42]. This mechanism performs global contextual modeling on features, thereby enhancing the model’s comprehension of inter-character relationships and the overall text image. Both the number of groups in the grouped convolution and the number of heads in MHSA are set to . Similar to SVTR, by adjusting hyper-parameters and , we can derive three variants of SVTRv2 with different capacities, i.e., Tiny, Small, and Base, which are detailed in Sec. 7 of Supplementary.
3.3 Feature Rearranging Module
We propose a feature rearrangement module (FRM) to tackle the CTC alignment puzzle arising from rotated text. It rearranges the 2D features F into a feature sequence synchronized with the reading order of the text image. We regard this process as mapping the relevant features from to , where and . This this rearrangement can be formalized by a matrix , whereby can be derived from .
Grounding in the fact that the degree of text curve and rotation can be decomposed into offset components in both horizontal and vertical directions, we propose to learn M by using two distinct steps: horizontal rearrangement and vertical rearrangement. As described in Eq. 1, the horizontal one operates on , each row in the feature map, to learn a horizontal rearrangement matrix , where element represents the probability that the horizontal rearranged feature F corresponds to the original feature Fi,m. Based on the learned , the features in each row are rearranged on the horizontal direction. Subsequently, through a residual and MLP processing, we obtain F.
where are learnable weights, is the Softmax function, and F.
In the following vertical rearrangement, we introduce a selecting token, denoted as , which simultaneously attends to each column of features F within Fh to learn a vertical rearrangement matrix , as detailed in Eq. 2. The element represents the probability that the vertical rearranged feature F corresponds to F. Moreover, all column features share a single selecting token, rather than assigning a unique selecting token to each column feature at different locations. This scheme allows the model to generalize to longer text sequences, even when the number of column features exceeds the number seen during training, thereby facilitating the effective recognition of long text.
where are learnable weights.
We denote F as the rearranged feature sequence . Then, the predicted character sequence is obtained after passes through the classifier , and further aligned with the label sequence using the CTC rule, where is the learnable weight of the classifier. Furthermore, to intuitively clarify the role of FRM, we rewrite the mapping relation between and the original feature F by the following procedure:
| M | ||||
As seen in Eq. 4, the visual features F, which fully describe the text to be recognized but may be irregularly organized, are mapped to a more canonical . Different from existing text rectification models [40, 32], which apply geometric transformations at image space and often fail to correct severely irregular text, our FRM is carried out at feature space. It learns a condensed feature sequence by mapping and rearranging the relevant cues from F to proper positions in . By employing a structured way that the features are first rearranged horizontally and then vertically, FRM can accommodate text instances with significant irregularities such as large orientation variations, and obtaining features better aligned with the CTC classification.
3.4 Semantic Guidance Module
CTC models directly classify visual features to obtain recognition results. This scheme inherently requires that the linguistic context must be incorporated into visual features, so that the CTC could benefit from. In light of this, we propose a semantic guidance module (SGM) as follows.
For each character in a text image with character labels Y = {, , , }, we define its contextual information as the surrounding left string = and right string = , where denotes the context window length. The SGM’s role is to guide the visual model to integrate context from both and into visual features.
We describe the process using the left string , with the same process applied symmetrically to the right string . Firstly, the characters in are mapped to string embeddings . Subsequently, these embeddings are encoded to create a hidden representation , representing the context of the left-side string . The attention map is computed by applying a dot product between the hidden representation and the visual features F, transformed by learned weight matrices and . The detailed formulation is as follows:
where represents a pre-defined token that encodes the left-side string. The attention map is used to weight the visual features F, producing the feature corresponding to character . After processing through the classifier , the predicted class probabilities for is obtained to calculate the cross-entropy loss with the label , where is learnable weights of the classifier.
The weight of the attention map records the relevance of to the visual features F, and moreover, represents the context of string . So only when the visual model incorporates the context from into the visual features of the target character , the attention map can maximize the relevance between and visual features of the character , thereby accurately highlighting the corresponding position of character , as shown in Fig. 3. A similar process applies to the right-side string , where the corresponding attention map and visual feature contribute to the prediction . By leveraging the above scheme during training, SGM effectively guides the visual model in integrating the linguistic context into visual features. Consequently, even when SGM is not used during inference, the linguistic context can still be maintained alongside the visual features, and enhancing the accuracy of CTC models. In contrast, previous methods, such as VisionLAN [48] and LPV [58], despite modeling linguistic context using visual features, still rely on attention-based decoders to unleash linguistic information, a process that is incompatible with CTC models.
3.5 Optimization Objective
During training, The optimization objective is to minimize the loss , which comprises and , as listed below:
where represents the cross-entropy loss, and are weighting parameters setting to 0.1 and 1, respectively.
4 Experiments
4.1 Datasets and Implementation Details
We evaluate SVTRv2 across multiple benchmarks covering diverse scenarios. They are: 1) six common regular and irregular benchmarks (Com), including ICDAR 2013 (IC13) [27], Street View Text (SVT) [45], IIIT5K-Words (IIIT5K) [33], ICDAR 2015 (IC15) [26], Street View Text-Perspective (SVTP) [35] and CUTE80 [1]. For IC13 and IC15, we use the versions with 857 and 1811 images, respectively; 2) the test set of the recent Union14M-L benchmark (U14M) [25], which includes seven challenging subsets: Curve, Multi-Oriented (MO), Artistic, Contextless (Cless), Salient, Multi-Words (MW) and General; 3) occluded scene text dataset (OST) [48], which is categorized into two subsets based on the degree of occlusion: weak occlusion (OSTw) and heavy occlusion (OSTh); 4) long text benchmark (LTB) [13], which includes 3376 samples of text length from 25 to 35; 5) the test set of BCTR [6], a Chinese text recognition benchmark with four subsets: Scene, Web, Document (Doc) and Hand-Writing (HW).
For English recognition, we train models on real-world datasets, from which the models exhibit stronger recognition capability [4, 25, 37]. There are three large-scale real-world training sets, i.e., the Real dataset [4], REBU-Syn [37], and the training set of Union14M-L (U14M-Train) [25]. However, they all overlap with U14M (detailed in Sec. 8 in Supplementary) across the seven subsets, leading to data leakage, which makes them unsuitable for training models. To resolve this, we introduce a filtered version of Union14M-L training set, termed as U14M-Filter, by filtering these overlapping instances. This new dataset is used to train SVTRv2 and 24 mainstream methods we reproduced.
For Chinese recognition, we train models on the training set of BCTR [6]. Unlike previous methods that train separately for each subset, we trained the model on an integrated dataset and then evaluated it on the four subsets.
We use AdamW optimizer [30] with a weight decay of 0.05 for training. The LR is set to and batchsize is set to 1024. One cycle LR scheduler [43] with 1.5/4.5 epochs linear warm-up is used in all the 20/100 epochs, where a/b means a for English and b for Chinese. For English model, we take SVTRv2 without SGM as the pre-trained and fine-tuned SVTRv2 with SGM with the same above settings. Word accuracy is used as the evaluation metric. Data augmentation like rotation, perspective distortion, motion blur and gaussian noise, are randomly performed and the maximum text length is set to 25 during training. The size of the character set is set to 94 for English and 6624 [28] for Chinese. In experiments, SVTRv2 means SVTRv2-B unless specified. All models are trained with mixed-precision on 4 RTX 4090 GPUs.
4.2 Ablation Study
Effectiveness of MSR. We group the Curve and MO text in U14M based on the aspect ratio Ri. As shown in Tab. 1, the majority of irregular texts fall within R1 and R2, where they are particularly prone to distortion when resized to a fixed size (see Fixed32×128 in Fig. 4). In contrast, MSR demonstrates significant improvements of 15.3% in R1 and 5.2% in R2 compared to Fixed32×128. Meanwhile, a large fixed-size Fixed64×256, although improving the accuracy compared to the baseline, still performs worse than our MSR by clear margins. The results strongly confirm our hypothesis that undesired resizing would hurt the recognition. Our MSR effectively mitigates this issue, providing better visual features thus enhancing the recognition accuracy.
Effectiveness of FRM. We ablate the two rearrangement sub-modules (Horizontal (H) rearranging and Vertical (V) rearranging). As shown in Tab. 1, compared to without FRM (w/o FRM), they individually improve accuracy by 2.03% and 0.71% on MO, and they together result in a 2.46% gain. Additionally, we explore using a Transformer block (+ TF1) to learn the rearrangement matrix holistically, whose effectiveness is less obvious. The most probable reason is that this scheme does not well distinguish between vertical and horizontal orientations. In contrast, FRM performs feature rearrangement in both directions, making it highly sensitive to text irregularity, and thus facilitating accurate CTC alignment. As shown in the left five cases in Fig. 4, FRM successfully recognizes reverse instances, providing strong evidence of FRM’s effectiveness.
Curve | MO | Com | U14M | ||||||||||||||||||||||||||||||||||||||
| SVTRv2 (+MSR+FRM) | 87.4 | 88.3 | 86.1 | 87.5 | 88.17 | 86.19 | 96.16 | 83.86 | |||||||||||||||||||||||||||||||||||||
| SVTRv2 (w/o both) | 70.5 | 81.5 | 82.8 | 84.4 | 82.89 | 65.59 | 95.28 | 77.78 | |||||||||||||||||||||||||||||||||||||
| vs. MSR (+FRM) | Fixed32×128 | 72.1 | 83.1 | 84.1 | 85.6 | 83.18 | 68.71 | 95.56 | 78.87 | ||||||||||||||||||||||||||||||||||||
| Padding32×W | 52.1 | 71.3 | 82.3 | 87.4 | 71.06 | 51.57 | 94.70 | 71.82 | |||||||||||||||||||||||||||||||||||||
| Fixed64×256 | 76.6 | 81.6 | 81.9 | 80.2 | 85.70 | 67.49 | 95.07 | 79.03 | |||||||||||||||||||||||||||||||||||||
| vs. FRM (+MSR) | w/o FRM | 85.7 | 86.3 | 86.0 | 85.5 | 87.35 | 83.73 | 95.44 | 82.22 | ||||||||||||||||||||||||||||||||||||
| + H rearranging | 87.0 | 87.1 | 86.3 | 85.5 | 88.05 | 85.76 | 95.98 | 82.94 | |||||||||||||||||||||||||||||||||||||
| + V rearranging | 85.0 | 87.6 | 88.5 | 85.5 | 88.01 | 84.44 | 95.66 | 82.70 | |||||||||||||||||||||||||||||||||||||
| + TF1 | 86.4 | 86.3 | 87.5 | 86.1 | 87.51 | 85.50 | 95.60 | 82.49 | |||||||||||||||||||||||||||||||||||||
| - | ResNet+TF3 | 49.3 | 63.5 | 64.0 | 66.7 | 65.00 | 42.07 | 92.26 | 63.00 | ||||||||||||||||||||||||||||||||||||
| FocalNet-B | 56.7 | 73.2 | 75.3 | 73.9 | 76.46 | 45.80 | 94.49 | 71.63 | |||||||||||||||||||||||||||||||||||||
| ConvNeXtV2 | 58.4 | 71.0 | 73.6 | 71.2 | 75.97 | 45.95 | 93.93 | 70.43 | |||||||||||||||||||||||||||||||||||||
| ViT-S | 68.5 | 73.8 | 73.8 | 73.0 | 75.02 | 64.35 | 93.57 | 72.09 | |||||||||||||||||||||||||||||||||||||
| SVTR-B | 53.3 | 74.8 | 76.4 | 78.4 | 76.22 | 44.49 | 94.58 | 71.17 | |||||||||||||||||||||||||||||||||||||
| +FRM | ResNet+TF3 | 53.8 | 67.9 | 65.5 | 65.8 | 69.00 | 46.02 | 93.12 | 66.81 | ||||||||||||||||||||||||||||||||||||
| FocalNet-B | 57.1 | 75.2 | 77.1 | 78.4 | 75.52 | 51.21 | 94.39 | 72.73 | |||||||||||||||||||||||||||||||||||||
| ConvNeXtV2 | 60.7 | 79.0 | 79.0 | 81.1 | 79.72 | 53.32 | 94.19 | 73.09 | |||||||||||||||||||||||||||||||||||||
| ViT-S | 75.1 | 79.4 | 79.0 | 78.4 | 80.42 | 72.17 | 94.44 | 77.07 | |||||||||||||||||||||||||||||||||||||
| SVTR-B | 59.1 | 79.0 | 78.8 | 80.2 | 79.84 | 51.28 | 94.75 | 73.48 | |||||||||||||||||||||||||||||||||||||
| +MSR | ResNet+TF3 | 68.2 | 71.3 | 75.3 | 72.1 | 75.64 | 60.33 | 93.50 | 71.95 | ||||||||||||||||||||||||||||||||||||
| FocalNet-B | 80.5 | 80.6 | 79.2 | 85.0 | 82.26 | 74.82 | 94.92 | 78.94 | |||||||||||||||||||||||||||||||||||||
| ConvNeXtV2 | 76.2 | 79.0 | 82.3 | 80.2 | 81.05 | 73.27 | 94.60 | 77.71 | |||||||||||||||||||||||||||||||||||||
| Method | OSTw | OSTh | Avg | Com | U14M | |
| Linguistic context modeling | w/o SGM | 82.86 | 66.97 | 74.92 | 96.16 | 83.86 |
| SGM | 86.26 | 73.80 | 80.03 | 96.57 | 86.14 | |
| GTC [23] | 83.07 | 68.32 | 75.70 | 96.01 | 84.33 | |
| ABINet [15] | 83.07 | 67.54 | 75.31 | 96.25 | 84.17 | |
| VisionLAN [48] | 83.25 | 68.97 | 76.11 | 96.39 | 84.01 | |
| PARSeq [4] | 83.85 | 69.24 | 76.55 | 96.21 | 84.72 | |
| MAERec [25] | 83.21 | 69.69 | 76.45 | 96.47 | 84.69 |
| Method | Venue | Encoder | Common Benchmarks | Avg | Union14M Benchmarks | Avg | LTB | OST | Size | FPS | ||||||||||||||||||||||||||
| ASTER [40] | TPAMI19 | ResNet+LSTM | 96.1 | 93.0 | 94.9 | 86.1 | 87.9 | 92.0 | 91.70 | 70.9 | 82.2 | 56.7 | 62.9 | 73.9 | 58.5 | 76.3 | 68.75 | 0.1 | 61.9 | 19.0 | 67.1 | |||||||||||||||
| NRTR [38] | ICDAR19 | Stem+TF6 | 98.1 | 96.8 | 97.8 | 88.9 | 93.3 | 94.4 | 94.89 | 67.9 | 42.4 | 66.5 | 73.6 | 66.4 | 77.2 | 78.3 | 67.46 | 0.0 | 74.8 | 44.3 | 17.3 | |||||||||||||||
| MORAN [32] | PR19 | ResNet+LSTM | 96.7 | 91.7 | 94.6 | 84.6 | 85.7 | 90.3 | 90.61 | 51.2 | 15.5 | 51.3 | 61.2 | 43.2 | 64.1 | 69.3 | 50.82 | 0.1 | 57.9 | 17.4 | 59.5 | |||||||||||||||
| SAR [29] | AAAI19 | ResNet+LSTM | 98.1 | 93.8 | 96.7 | 86.0 | 87.9 | 95.5 | 93.01 | 70.5 | 51.8 | 63.7 | 73.9 | 64.0 | 79.1 | 75.5 | 68.36 | 0.0 | 60.6 | 57.5 | 15.8 | |||||||||||||||
| DAN [47] | AAAI20 | ResNet+FPN | 97.5 | 94.7 | 96.5 | 87.1 | 89.1 | 94.4 | 93.24 | 74.9 | 63.3 | 63.4 | 70.6 | 70.2 | 71.1 | 76.8 | 70.05 | 0.0 | 61.8 | 27.7 | 99.0 | |||||||||||||||
| SRN [55] | CVPR20 | ResNet+FPN | 97.2 | 96.3 | 97.5 | 87.9 | 90.9 | 96.9 | 94.45 | 78.1 | 63.2 | 66.3 | 65.3 | 71.4 | 58.3 | 76.5 | 68.43 | 0.0 | 64.6 | 51.7 | 67.1 | |||||||||||||||
| SEED [36] | CVPR20 | ResNet+LSTM | 96.5 | 93.2 | 94.2 | 87.5 | 88.7 | 93.4 | 92.24 | 69.1 | 80.9 | 56.9 | 63.9 | 73.4 | 61.3 | 76.5 | 68.87 | 0.1 | 62.6 | 24.0 | 65.4 | |||||||||||||||
| AutoSTR [59] | ECCV20 | NAS+LSTM | 96.8 | 92.4 | 95.7 | 86.6 | 88.2 | 93.4 | 92.19 | 72.1 | 81.7 | 56.7 | 64.8 | 75.4 | 64.0 | 75.9 | 70.09 | 0.1 | 61.5 | 6.0 | 82.6 | |||||||||||||||
| RoScanner [57] | ECCV20 | ResNet | 98.5 | 95.8 | 97.7 | 88.2 | 90.1 | 97.6 | 94.65 | 79.4 | 68.1 | 70.5 | 79.6 | 71.6 | 82.5 | 80.8 | 76.08 | 0.0 | 68.6 | 48.0 | 64.1 | |||||||||||||||
| ABINet [15] | CVPR21 | ResNet+TF3 | 98.5 | 98.1 | 97.7 | 90.1 | 94.1 | 96.5 | 95.83 | 80.4 | 69.0 | 71.7 | 74.7 | 77.6 | 76.8 | 79.8 | 75.72 | 0.0 | 75.0 | 36.9 | 73.0 | |||||||||||||||
| VisionLAN [48] | ICCV21 | ResNet+TF3 | 98.2 | 95.8 | 97.1 | 88.6 | 91.2 | 96.2 | 94.50 | 79.6 | 71.4 | 67.9 | 73.7 | 76.1 | 73.9 | 79.1 | 74.53 | 0.0 | 66.4 | 32.9 | 93.5 | |||||||||||||||
| PARSeq [4] | ECCV22 | ViT-S | 98.9 | 98.1 | 98.4 | 90.1 | 94.3 | 98.6 | 96.40 | 87.6 | 88.8 | 76.5 | 83.4 | 84.4 | 84.3 | 84.9 | 84.26 | 0.0 | 79.9 | 23.8 | 52.6 | |||||||||||||||
| MATRN [34] | ECCV22 | ResNet+TF3 | 98.8 | 98.3 | 97.9 | 90.3 | 95.2 | 97.2 | 96.29 | 82.2 | 73.0 | 73.4 | 76.9 | 79.4 | 77.4 | 81.0 | 77.62 | 0.0 | 77.8 | 44.3 | 46.9 | |||||||||||||||
| MGP-STR [46] | ECCV22 | ViT-B | 97.9 | 97.8 | 97.1 | 89.6 | 95.2 | 96.9 | 95.75 | 85.2 | 83.7 | 72.6 | 75.1 | 79.8 | 71.1 | 83.1 | 78.65 | 0.0 | 78.7 | 148 | 120 | |||||||||||||||
| CPPD [12] | Preprint | SVTR-B | 99.0 | 97.8 | 98.2 | 90.4 | 94.0 | 99.0 | 96.40 | 86.2 | 78.7 | 76.5 | 82.9 | 83.5 | 81.9 | 83.5 | 81.91 | 0.0 | 79.6 | 27.0 | 125 | |||||||||||||||
| LPV [58] | IJCAI23 | SVTR-B | 98.6 | 97.8 | 98.1 | 89.8 | 93.6 | 97.6 | 95.93 | 86.2 | 78.7 | 75.8 | 80.2 | 82.9 | 81.6 | 82.9 | 81.20 | 0.0 | 77.7 | 30.5 | 82.6 | |||||||||||||||
| MAERec [25] | ICCV23 | ViT-S | 99.2 | 97.8 | 98.2 | 90.4 | 94.3 | 98.3 | 96.36 | 89.1 | 87.1 | 79.0 | 84.2 | 86.3 | 85.9 | 84.6 | 85.17 | 9.8 | 76.4 | 35.7 | 17.1 | |||||||||||||||
| LISTER [8] | ICCV23 | FocalNet-B | 98.8 | 97.5 | 98.6 | 90.0 | 94.4 | 96.9 | 96.03 | 78.7 | 68.8 | 73.7 | 81.6 | 74.8 | 82.4 | 83.5 | 77.64 | 36.3 | 77.1 | 51.1 | 44.6 | |||||||||||||||
| CDistNet [62] | IJCV24 | ResNet+TF3 | 98.7 | 97.1 | 97.8 | 89.6 | 93.5 | 96.9 | 95.59 | 81.7 | 77.1 | 72.6 | 78.2 | 79.9 | 79.7 | 81.1 | 78.62 | 0.0 | 71.8 | 43.3 | 15.9 | |||||||||||||||
| CAM [54] | PR24 | ConvNeXtV2 | 98.2 | 96.1 | 96.6 | 89.0 | 93.5 | 96.2 | 94.94 | 85.4 | 89.0 | 72.0 | 75.4 | 84.0 | 74.8 | 83.1 | 80.52 | 0.7 | 74.2 | 58.7 | 28.6 | |||||||||||||||
| BUSNet [49] | AAAI24 | ViT-S | 98.3 | 98.1 | 97.8 | 90.2 | 95.3 | 96.5 | 96.06 | 83.0 | 82.3 | 70.8 | 77.9 | 78.8 | 71.2 | 82.6 | 78.10 | 0.0 | 78.7 | 32.1 | 83.3 | |||||||||||||||
| OTE [52] | CVPR24 | SVTR-B | 98.6 | 96.6 | 98.0 | 90.1 | 94.0 | 97.2 | 95.74 | 86.0 | 75.8 | 74.6 | 74.7 | 81.0 | 65.3 | 82.3 | 77.09 | 0.0 | 77.8 | 20.3 | 55.2 | |||||||||||||||
| CRNN [39] | TPAMI16 | ResNet+LSTM | 95.8 | 91.8 | 94.6 | 84.9 | 83.1 | 91.0 | 90.21 | 48.1 | 13.0 | 51.2 | 62.3 | 41.4 | 60.4 | 68.2 | 49.24 | 47.2 | 58.0 | 16.2 | 172 | |||||||||||||||
| SVTR [11] | IJCAI22 | SVTR-B | 98.0 | 97.1 | 97.3 | 88.6 | 90.7 | 95.8 | 94.58 | 76.2 | 44.5 | 67.8 | 78.7 | 75.2 | 77.9 | 77.8 | 71.17 | 45.1 | 69.6 | 18.1 | 161 | |||||||||||||||
| SVTRv2-T | 98.6 | 96.6 | 98.0 | 88.4 | 90.5 | 96.5 | 94.78 | 83.6 | 76.0 | 71.2 | 82.4 | 77.2 | 82.3 | 80.7 | 79.05 | 47.8 | 71.4 | 5.1 | 201 | |||||||||||||||||
| SVTRv2-S | 99.0 | 98.3 | 98.5 | 89.5 | 92.9 | 98.6 | 96.13 | 88.3 | 84.6 | 76.5 | 84.3 | 83.3 | 85.4 | 83.5 | 83.70 | 47.6 | 78.0 | 11.3 | 189 | |||||||||||||||||
| C T C | SVTRv2 | - | SVTRv2-B | 99.2 | 98.0 | 98.7 | 91.1 | 93.5 | 99.0 | 96.57 | 90.6 | 89.0 | 79.3 | 86.1 | 86.2 | 86.7 | 85.1 | 86.14 | 50.2 | 80.0 | 19.8 | 143 | ||||||||||||||
Effectiveness of SGM. As illustrated in Tab. 2, SGM achieves 0.41% and 2.28% increase on Com and U14M, respectively, while gains a 5.11% improvement on OST. Since OST frequently suffers from missing a portion of characters, this notable gain implies that the linguistic context has been successfully established. For comparison, we also employ GTC [23] and four popular language decoders [15, 48, 4, 25] to substitute for our SGM. However, there is no much difference between the gains obtained from OST and the other two datasets (Com and U14M). This suggests that SGM offers a distinct advantage in integrating linguistic context into visual features, and significantly improving the recognition accuracy of CTC models. The five cases on the right side of Fig. 4 showcase that SGM facilitates SVTRv2 to accurately decipher occluded characters, achieving comparable results with PARSeq [4], which is equipped with an advanced permuted language model.
Adaptability to different visual models. We further examine MSR, FRM, and SGM on five frequently used visual models [20, 10, 11, 53, 50]. As presented in the bottom part of Tab. 1, these modules consistently enhance the performance (ViT [10] and SVTR [11] employ absolute positional coding and do not compatible with MSR). When both FRM and MSR modules incorporated, ResNet+TF3 [20], FocalNet [53], and ConvNeXtV2 [54] exhibit significant accuracy improvements, either matching or even exceeding the accuracy of their EDTR counterparts (see Tab. 3). The results highlight the versatility of the three proposed modules.
4.3 Comparison with State-of-the-arts
To demonstrate the effectiveness of SVTRv2 in English, we compare it with 24 popular STR methods. All the models are tested on the newly constructed U14M and the results are given in Tab. 3. SVTRv2-B ranks the top in 12 of the 15 evaluated scenarios, It almost outperforms all the EDTRs in every scenario, showing a clear accuracy advantage. Meanwhile, it still enjoys a small model size and a significant speed advantage. Specifically, compared to MAERec, the best-performed existing model on U14M, SVTRv2-B shows an accuracy improvement of 0.97% and 8 faster inference speed. Compared to CPPD, which is known for its accuracy-speed tradeoff, SVTR-B runs faster than 10%, along with a 4.23% accuracy increase on U14M. Regarding OST, as illustrated in the right part of Fig. 4, SVTR-B relies solely on a single visual model but achieves comparable accuracy to PARSeq, which employed the advanced permuted language model and is the best-performed existing model on OST. In the case of long text recognition, where a large portion of EDTRs are incapable of recognizing, SVTR-B outperforms LISTER, the best EDTR method on LTB, by 13%, demonstrating the remarkable scalability of SVTRv2. In addition, SVTRv2-T and SVTRv2-S, the two smaller models also show leading accuracy compared with models of similar sizes, offering solutions with different accuracy-speed tradeoff.
Two observations are derived when looking into the results on Curve and MO. First, SVTRv2 models significantly surpass existing CTC models. For example, compared to SVTR-B, SVTRv2-B gains prominent accuracy improvements of 14.4% and 44.5%, respectively. Second, as shown in Tab. 4, comparing with previous methods employing the rectification modules [40, 36, 59, 62, 54, 11, 61] and the attention-based decoder [38, 31, 5, 51, 56, 25, 54, 52, 47, 29] to recognize irregular text, SVTRv2 also performs better than these methods on Curve and MO. In Fig. 4, the rectification module (TPS) and the attention-based decoder (MAERec*) do not recognize the extremely curved and rotated text correctly, in contrast, SMTR successes. Moreover, as demonstrated by the results on LTB in Tab. 4 and Fig. 5, TPS and MAERec* both do not effectively recognize long text, while SVTRv2 circumvents this limitation. These results indicate that our proposed modules successfully address the challenge of handling irregular text that existing CTC models encountered, while preserving CTC’s proficiency in recognizing long text.
SVTRv2 models also exhibit strong performance in Chinese text recognition (see Tab. 5), where SVTRv2-B achieve state of the art. The result underscores its great adaptability to different languages. To sum, we evaluate SVTRv2 across a wide range of scenarios. The results consistently confirm that this CTC model beats leading EDTRs.
| R1 | R2 | R3 | R4 | Curve | MO | Com | U14M | LTB | ||
| SVTRv2 | 90.8 | 89.0 | 90.4 | 91.0 | 90.64 | 89.04 | 96.57 | 86.14 | 50.2 | |
| TPS | SVTR [11] | 86.8 | 82.3 | 77.3 | 75.7 | 82.19 | 86.12 | 94.62 | 78.44 | 0.0 |
| SVTRv2 | 89.5 | 85.1 | 78.4 | 83.8 | 84.71 | 88.97 | 94.62 | 79.94 | 0.5 | |
| MAE- REC* | SVTR [11] | 81.3 | 87.6 | 87.6 | 88.3 | 87.88 | 78.74 | 96.32 | 83.23 | 0.0 |
| SVTRv2 | 88.0 | 88.9 | 89.4 | 88.3 | 89.96 | 87.56 | 96.42 | 85.67 | 0.2 | |
| Method | Scene | Web | Doc | HW | Avg | SceneL>25 | Size |
| ASTER [40] | 61.3 | 51.7 | 96.2 | 37.0 | 61.55 | - | 27.2 |
| MORAN [32] | 54.6 | 31.5 | 86.1 | 16.2 | 47.10 | - | 28.5 |
| SAR [29] | 59.7 | 58.0 | 95.7 | 36.5 | 62.48 | - | 27.8 |
| SEED [36] | 44.7 | 28.1 | 91.4 | 21.0 | 46.30 | - | 36.1 |
| MASTER [31] | 62.8 | 52.1 | 84.4 | 26.9 | 56.55 | - | 62.8 |
| ABINet [15] | 66.6 | 63.2 | 98.2 | 53.1 | 70.28 | - | 53.1 |
| TransOCR [5] | 71.3 | 64.8 | 97.1 | 53.0 | 71.55 | - | 83.9 |
| CCR-CLIP [56] | 71.3 | 69.2 | 98.3 | 60.3 | 74.78 | - | 62.0 |
| DCTC [60] | 73.9 | 68.5 | 99.4 | 51.0 | 73.20 | - | 40.8 |
| CAM [54] | 76.0 | 69.3 | 98.1 | 59.2 | 76.80 | - | 135 |
| PARSeq* [4] | 84.2 | 82.8 | 99.5 | 63.0 | 82.37 | 0.0 | 28.9 |
| CPPD* [12] | 82.7 | 82.4 | 99.4 | 62.3 | 81.72 | 0.0 | 32.1 |
| MAERec* [25] | 84.4 | 83.0 | 99.5 | 65.6 | 83.13 | 4.1 | 40.8 |
| LISTER* [8] | 79.4 | 79.5 | 99.2 | 58.0 | 79.02 | 13.9 | 55.0 |
| CRNN* [39] | 63.8 | 68.2 | 97.0 | 46.1 | 68.76 | 37.6 | 19.5 |
| SVTR-B* [11] | 77.9 | 78.7 | 99.2 | 62.1 | 79.49 | 22.9 | 19.8 |
| SVTRv2-T | 77.8 | 78.8 | 99.2 | 62.0 | 79.45 | 47.8 | 6.8 |
| SVTRv2-S | 81.1 | 81.2 | 99.3 | 65.0 | 81.64 | 50.0 | 14.0 |
| SVTRv2-B | 83.5 | 83.3 | 99.5 | 67.0 | 83.31 | 52.8 | 22.5 |
5 Conclusion
In this paper, we have presented SVTRv2, an accurate and efficient CTC-based STR method. SVTRv2 is featured by developing the MSR and FRM modules to tackle the text irregular challenge, and devising the SGM module to endow linguistic context to the visual model. These upgrades maintain the simple architecture of CTC models, thus they remain quite efficient. More importantly, our thorough validation on multiple benchmarks demonstrates the effectiveness of SVTRv2. It achieves leading accuracy and inference speed in various challenging scenarios covering regular, irregular, occluded, Chinese and long text, convincingly indicating that SVTRv2 has beat EDTRs in scene text recognition. In addition, we retrain 24 methods from scratch on U14M-Filter without data leakage, constituting a comprehensive and reliable benchmark. We hope that SVTRv2 and this benchmark will further advance the development of the OCR community.
Supplementary Material
| ID | Method | Common Benchmarks | Avg | Union14M Benchmarks | Avg | LTB | OST | Size | FPS | |||||||||||||||||||||||||
| 0 | SVTR (w/ TPS) | 98.1 | 96.1 | 96.4 | 89.2 | 92.1 | 95.8 | 94.62 | 82.2 | 86.1 | 69.7 | 75.1 | 81.6 | 73.8 | 80.7 | 78.44 | 0.0 | 71.2 | 19.95 | 141 | ||||||||||||||
| 1 | 0 + w/o TPS | 98.0 | 97.1 | 97.3 | 88.6 | 90.7 | 95.8 | 94.58 | 76.2 | 44.5 | 67.8 | 78.7 | 75.2 | 77.9 | 77.8 | 71.17 | 45.1 | 67.8 | 18.10 | 161 | ||||||||||||||
| 2 | 1 + | 98.9 | 97.4 | 97.9 | 89.7 | 91.8 | 96.9 | 95.41 | 82.2 | 64.3 | 70.2 | 80.0 | 80.9 | 80.6 | 80.5 | 76.95 | 44.8 | 69.5 | 18.10 | 145 | ||||||||||||||
| 3 | 2 + Conv2 | 98.7 | 97.1 | 97.1 | 89.6 | 91.6 | 97.6 | 95.28 | 82.9 | 65.6 | 73.2 | 80.0 | 80.5 | 81.6 | 80.8 | 77.78 | 47.4 | 71.1 | 17.77 | 159 | ||||||||||||||
| 4 | 3 + MSR | 98.7 | 98.0 | 97.4 | 89.4 | 91.6 | 97.6 | 95.44 | 87.4 | 83.7 | 75.4 | 80.9 | 81.9 | 83.5 | 82.8 | 82.22 | 50.9 | 72.5 | 17.77 | 159 | ||||||||||||||
| 5 | 4 + FRM | 98.8 | 98.1 | 98.4 | 89.8 | 92.9 | 99.0 | 96.16 | 88.2 | 86.2 | 77.5 | 83.2 | 83.9 | 84.6 | 83.5 | 83.86 | 50.7 | 74.9 | 19.76 | 143 | ||||||||||||||
| 6 | 5 + SGM | 99.2 | 98.0 | 98.7 | 91.1 | 93.5 | 99.0 | 96.57 | 90.6 | 89.0 | 79.3 | 86.1 | 86.2 | 86.7 | 85.1 | 86.14 | 50.2 | 80.0 | 19.76 | 143 | ||||||||||||||
6 More detail of Ablation Study
SVTRv2 builds upon the foundation of SVTR by introducing several innovative strategies aimed at addressing challenges in recognizing irregular text and modeling linguistic context. The key advancements and their impact are systematically detailed below:
Removal of the Rectification Module and Introduction of MSR and FRM. In the original SVTR, a rectification module is employed to recognize irregular texts. However, this approach negatively impacts the recognition of long texts. To overcome this limitation, SVTRv2 removes the rectification module entirely. To effectively handle irregular text without compromising the CTC model’s ability to generalize to long text, MSR and FRM are introduced.
Improvement in Feature Resolution. SVTR extracts visual representations of size from input images of size . While this approach is effective for regular text, it struggles with retaining the distinct characteristics of irregular text. SVTRv2 doubles the height resolution () of visual features, producing features of size , thereby improving its capacity to recognize irregular text.
Refinement of Local Mixing Mechanisms. SVTR employs a hierarchical vision transformer structure, leveraging two mixing strategies: Local Mixing is implemented through a sliding window-based local attention mechanism, and Global Mixing employs the standard global multi-head self-attention mechanism. SVTRv2 retains the hierarchical vision transformer structure and the global multi-head self-attention mechanism for Global Mixing. For Local Mixing, SVTRv2 introduces a pivotal change. Specifically, the sliding window-based local attention is replaced with two consecutive group convolutions (Conv2). It is important to highlight that unlike previous CNNs [41, 20, 22], there is no normalization or activation layer between the two convolutions.
Semantic Guidance Module (SGM). The original SVTR model relies solely on CTC framework for both training and inference. However, CTC is inherently limited in its ability to model linguistic context. SVTRv2 addresses this by introducing a Semantic Guidance Module (SGM) during training. SGM facilitates the visual encoder in capturing linguistic information, enriching the feature representation. Importantly, SGM is discarded during inference, ensuring that the efficiency of CTC-based decoding remains unaffected while still benefiting from its contributions during the training phase.
6.1 Progressive Ablation Experiments
To comprehensively evaluate the contributions of the innovations in SVTRv2, a series of progressive ablation experiments are conducted. Tab 6 outlines the results, with the following observations:
1. Baseline (ID0): The original SVTR serves as the baseline for comparison.
2. Rectification Module Removal (ID1) reveals that while the rectification module (e.g., TPS) improves irregular text recognition accuracy, it hinders the model’s ability to recognize long texts. This confirms its limitations in balancing these tasks.
3. Improvement in Feature Resolution (ID2): Doubling the height resolution () significantly boosts performance across challenging datasets, particularly for irregular text.
4. Replacement of Local Attention with Conv2 (ID3): Replacing the sliding window-based local attention with two consecutive group convolutions (Conv2) yields improvements in artistic text, with a 3.0% increase in accuracy. This result highlights the efficacy of convolution-based approaches in capturing character-level nuances, such as strokes and textures, thereby improving its ability to recognize artistic and irregular text styles.
5. Incorporation of MSR and FRM (ID4 and ID5): These components collectively enhance accuracy on irregular text benchmarks (e.g., Curve), surpassing the rectification-based SVTR (ID0) by 6.0%, without compromising the CTC model’s ability to generalize to long text.
6. Integration of SGM (ID6): Adding SGM yields significant gains on multiple datasets, improving accuracy on OST by 5.11% and Union14M-Benchmark by 2.28%.
It can be summarized as that, by integrating Conv2, MSR, FRM, and SGM, SVTRv2 significantly improves performance in recognizing irregular text and modeling linguistic context over SVTR, while maintaining robust long-text recognition capabilities and preserving the efficiency of CTC-based inference.
7 SVTRv2 Variants
There are several hyper-parameters in SVTRv2, including the depth of channel () and the number of heads at each stage, the number of mixing blocks () and their permutation. By varying them, SVTRv2 architectures with different capacities could be obtained and we construct three typical ones, i.e., SVTRv2-T (Tiny), SVTRv2-S (Small), SVTRv2-B (Base). Their detail configurations are shown in Tab. 7.
denotes that the first m mixing blocks in SVTRv2 utilize local mixing, while the last n mixing blocks employ global mixing. Specifically, in SVTRv2-T and SVTR-S, all blocks in the first stage and the first three blocks in the second stage use local mixing. The last three blocks in the second stage, as well as all blocks in the third stage, are global mixing. In the case of SVTRv2-B, all blocks in the first stage and the first two blocks in the second stage use local mixing, whereas the last four blocks in the second stage and all blocks in the third stage adopt global mixing.
| Models | Heads | Permutation | ||
| SVTRv2-T | [64,128,256] | [3,6,3] | [2,4,8] | |
| SVTRv2-S | [96,192,384] | [3,6,3] | [3,6,12] | |
| SVTRv2-B | [128,256,384] | [6,6,6] | [4,8,12] |
| Curve | MO | Artistic | Cless | Salient | MW | General | |
| 2,426 | 1,369 | 900 | 779 | 1,585 | 829 | 400,000 | |
| Real [4] | 1,276 | 440 | 432 | 326 | 431 | 193 | 254,174 |
| REBU-Syn [37] | 1,285 | 443 | 462 | 363 | 442 | 289 | 260,575 |
| U14M-Train [25] | 9 | 3 | 30 | 37 | 11 | 96 | 6,401 |
| Method | Venue | Encoder | Common Benchmarks | Avg | Union14M Benchmarks | Avg | Size | |||||||||||||||||||||||||
| ASTER [40] | TPAMI2019 | ResNet+LSTM | 93.3 | 90.0 | 90.8 | 74.7 | 80.2 | 80.9 | 84.98 | 34.0 | 10.2 | 27.7 | 33.0 | 48.2 | 27.6 | 39.8 | 31.50 | 27.2 | ||||||||||||||
| NRTR [38] | ICDAR2019 | Stem+TF6 | 90.1 | 91.5 | 95.8 | 79.4 | 86.6 | 80.9 | 87.38 | 31.7 | 4.40 | 36.6 | 37.3 | 30.6 | 54.9 | 48.0 | 34.79 | 31.7 | ||||||||||||||
| MORAN [32] | PR2019 | ResNet+LSTM | 91.0 | 83.9 | 91.3 | 68.4 | 73.3 | 75.7 | 80.60 | 8.90 | 0.70 | 29.4 | 20.7 | 17.9 | 23.8 | 35.2 | 19.51 | 17.4 | ||||||||||||||
| SAR [29] | AAAI2019 | ResNet+LSTM | 91.5 | 84.5 | 91.0 | 69.2 | 76.4 | 83.5 | 82.68 | 44.3 | 7.70 | 42.6 | 44.2 | 44.0 | 51.2 | 50.5 | 40.64 | 57.7 | ||||||||||||||
| DAN [47] | AAAI2020 | ResNet+FPN | 93.4 | 87.5 | 92.1 | 71.6 | 78.0 | 81.3 | 83.98 | 26.7 | 1.50 | 35.0 | 40.3 | 36.5 | 42.2 | 42.1 | 32.04 | 27.7 | ||||||||||||||
| SRN [55] | CVPR2020 | ResNet+FPN | 94.8 | 91.5 | 95.5 | 82.7 | 85.1 | 87.8 | 89.57 | 63.4 | 25.3 | 34.1 | 28.7 | 56.5 | 26.7 | 46.3 | 40.14 | 54.7 | ||||||||||||||
| SEED* [36] | CVPR2020 | ResNet+LSTM | 93.8 | 89.6 | 92.8 | 80.0 | 81.4 | 83.6 | 86.87 | 40.4 | 15.5 | 32.1 | 32.5 | 54.8 | 35.6 | 39.0 | 35.70 | 24.0 | ||||||||||||||
| AutoSTR* [59] | ECCV2020 | NAS+LSTM | 94.7 | 90.9 | 94.2 | 81.8 | 81.7 | - | - | 47.7 | 17.9 | 30.8 | 36.2 | 64.2 | 38.7 | 41.3 | 39.54 | 6.00 | ||||||||||||||
| RoScanner [57] | ECCV2020 | ResNet | 95.3 | 88.1 | 94.8 | 77.1 | 79.5 | 90.3 | 87.52 | 43.6 | 7.90 | 41.2 | 42.6 | 44.9 | 46.9 | 39.5 | 38.09 | 48.0 | ||||||||||||||
| ABINet [15] | CVPR2021 | ResNet+TF3 | 96.2 | 93.5 | 97.4 | 86.0 | 89.3 | 89.2 | 91.93 | 59.5 | 12.7 | 43.3 | 38.3 | 62.0 | 50.8 | 55.6 | 46.03 | 36.7 | ||||||||||||||
| VisionLAN [48] | ICCV2021 | ResNet+TF3 | 95.8 | 91.7 | 95.7 | 83.7 | 86.0 | 88.5 | 90.23 | 57.7 | 14.2 | 47.8 | 48.0 | 64.0 | 47.9 | 52.1 | 47.39 | 32.8 | ||||||||||||||
| PARSeq* [4] | ECCV2022 | ViT-S | 97.0 | 93.6 | 97.0 | 86.5 | 88.9 | 92.2 | 92.53 | 63.9 | 16.7 | 52.5 | 54.3 | 68.2 | 55.9 | 56.9 | 52.62 | 23.8 | ||||||||||||||
| MATRN [34] | ECCV2022 | ResNet+TF3 | 96.6 | 95.0 | 97.9 | 86.6 | 90.6 | 93.5 | 93.37 | 63.1 | 13.4 | 43.8 | 41.9 | 66.4 | 53.2 | 57.0 | 48.40 | 44.2 | ||||||||||||||
| MGP-STR* [46] | ECCV2022 | ViT-B | 96.4 | 94.7 | 97.3 | 87.2 | 91.0 | 90.3 | 92.82 | 55.2 | 14.0 | 52.8 | 48.5 | 65.2 | 48.8 | 59.1 | 49.09 | 148 | ||||||||||||||
| LevOCR* [9] | ECCV2022 | ResNet+TF3 | 96.6 | 94.4 | 96.7 | 86.5 | 88.8 | 90.6 | 92.27 | 52.8 | 10.7 | 44.8 | 51.9 | 61.3 | 54.0 | 58.1 | 47.66 | 109 | ||||||||||||||
| CornerTF* [51] | ECCV2022 | CornerEncoder | 95.9 | 94.6 | 97.8 | 86.5 | 91.5 | 92.0 | 93.05 | 62.9 | 18.6 | 56.1 | 58.5 | 68.6 | 59.7 | 61.0 | 55.07 | 86.0 | ||||||||||||||
| CPPD [12] | Preprint | SVTR-B | 97.6 | 95.5 | 98.2 | 87.9 | 90.9 | 92.7 | 93.80 | 65.5 | 18.6 | 56.0 | 61.9 | 71.0 | 57.5 | 65.8 | 56.63 | 26.8 | ||||||||||||||
| SIGA* [17] | CVPR2023 | ViT-B | 96.6 | 95.1 | 97.8 | 86.6 | 90.5 | 93.1 | 93.28 | 59.9 | 22.3 | 49.0 | 50.8 | 66.4 | 58.4 | 56.2 | 51.85 | 113 | ||||||||||||||
| CCD* [18] | ICCV2023 | ViT-B | 97.2 | 94.4 | 97.0 | 87.6 | 91.8 | 93.3 | 93.55 | 66.6 | 24.2 | 63.9 | 64.8 | 74.8 | 62.4 | 64.0 | 60.10 | 52.0 | ||||||||||||||
| LISTER* [8] | ICCV2023 | FocalNet-B | 96.9 | 93.8 | 97.9 | 87.5 | 89.6 | 90.6 | 92.72 | 56.5 | 17.2 | 52.8 | 63.5 | 63.2 | 59.6 | 65.4 | 54.05 | 49.9 | ||||||||||||||
| LPV-B* [58] | IJCAI2023 | SVTR-B | 97.3 | 94.6 | 97.6 | 87.5 | 90.9 | 94.8 | 93.78 | 68.3 | 21.0 | 59.6 | 65.1 | 76.2 | 63.6 | 62.0 | 59.40 | 35.1 | ||||||||||||||
| CDistNet* [62] | IJCV2024 | ResNet+TF3 | 96.4 | 93.5 | 97.4 | 86.0 | 88.7 | 93.4 | 92.57 | 69.3 | 24.4 | 49.8 | 55.6 | 72.8 | 64.3 | 58.5 | 56.38 | 65.5 | ||||||||||||||
| CAM* [54] | PR2024 | ConvNeXtV2-B | 97.4 | 96.1 | 97.2 | 87.8 | 90.6 | 92.4 | 93.58 | 63.1 | 19.4 | 55.4 | 58.5 | 72.7 | 51.4 | 57.4 | 53.99 | 135 | ||||||||||||||
| BUSNet [49] | AAAI2024 | ViT-S | 96.2 | 95.5 | 98.3 | 87.2 | 91.8 | 91.3 | 93.38 | - | - | - | - | - | - | - | - | 56.8 | ||||||||||||||
| DCTC [60] | AAAI2024 | SVTR-L | 96.9 | 93.7 | 97.4 | 87.3 | 88.5 | 92.3 | 92.68 | - | - | - | - | - | - | - | - | 40.8 | ||||||||||||||
| OTE [52] | CVPR2024 | SVTR-B | 96.4 | 95.5 | 97.4 | 87.2 | 89.6 | 92.4 | 93.08 | - | - | - | - | - | - | - | - | 25.2 | ||||||||||||||
| CRNN [39] | TPAMI2016 | ResNet+LSTM | 82.9 | 81.6 | 91.1 | 69.4 | 70.0 | 65.5 | 76.75 | 7.50 | 0.90 | 20.7 | 25.6 | 13.9 | 25.6 | 32.0 | 18.03 | 8.30 | ||||||||||||||
| SVTR* [11] | IJCAI2022 | SVTR-B | 96.0 | 91.5 | 97.1 | 85.2 | 89.9 | 91.7 | 91.90 | 69.8 | 37.7 | 47.9 | 61.4 | 66.8 | 44.8 | 61.0 | 55.63 | 24.6 | ||||||||||||||
| SVTRv2 | - | SVTRv2-B | 97.7 | 94.0 | 97.3 | 88.1 | 91.2 | 95.8 | 94.02 | 74.6 | 25.2 | 57.6 | 69.7 | 77.9 | 68.0 | 66.9 | 62.83 | 19.8 | ||||||||||||||
8 More detail of real-world datasets
For English recognition, we train models on real-world datasets, from which the models exhibit stronger recognition capability [4, 25, 37]. There are three large-scale real-world training sets, i.e., the Real dataset [4], REBU-Syn [37], and the training set of Union14M-L (U14M-Train) [25]. However, as shown in Fig. 6 and Tab. 8, the former two significantly overlap with Union14M-Benchmarks, thus not suitable for model training. Surprisingly, U14M-Train is also overlapped with Union14M-Benchmarks, in nearly 6.5k text instances across the seven subsets. It means the models trained based on U14M-Train suffer from data leakage when tested on Union14M-Benchmarks,, thus the results reported by [25] should be updated. To this end, we create a filtered version of Union14M-L training set, termed as U14M-Filter, by filtering these overlapping instances. This new dataset is used to train SVTRv2 and other 24 methods we reproduced.
9 More detail of Inference Time
In term of the inference time, we do not utilize any acceleration framework and instead employ PyTorch’s dynamic graph mode on one NVIDIA 1080Ti GPU. We first measure the inference time for 3,000 images with a batch size of 1, calculating the average inference time for each text length. We then computed the arithmetic mean of the average time across all text lengths to determine the overall inference time of the model. Algorithm 1 details the process of measuring inference time.
10 Results when trained on synthetic datasets
Previous research typically follows a typical evaluation protocol, where models are trained on synthetic datasets and validated using six widely recognized real-world benchmarks. Building on this protocol, we trained the SVTRv2 model on synthetic datasets. In addition to evaluating SVTRv2 on the six common benchmarks, we also assess its performance on challenging benchmarks, i.e. Union14M-Benchmark, offering a comprehensive understanding of the model’s generalization capabilities. For methods that have not reported performance on challenging benchmarks, we conducted additional evaluations using their publicly available models and present these results for comparative analysis. As illustrated in Tab. 9, models trained on synthetic datasets exhibit notably reduced performance compared to those trained on large-scale real-world datasets (see Tab. 3). This performance drop is particularly pronounced on challenging benchmarks. These findings emphasize the critical importance of real datasets in improving recognition accuracy for challenging text scenarios.
Despite the challenges associated with synthetic datasets, SVTRv2 exhibits superior performance across both average accuracy metrics, surpassing the previously best-performing method by 0.22% and 2.73%, respectively. On irregular text benchmarks, such as Curve and Multi-Oriented, SVTR achieves strong results, largely due to its integrated correction module, which is particularly adept at handling irregular text patterns, even when trained on synthetic datasets. Notably, SVTRv2 achieves a substantial 4.8% improvement over SVTR on the Curve benchmark, further demonstrating its enhanced capacity to address irregular text. Overall, these results demonstrate that, even when trained solely on synthetic datasets, SVTRv2 exhibits strong generalization capabilities, effectively handling complex and challenging text recognition scenarios.
11 Qualitative Analysis of Recognition Results
The SVTRv2 model achieved an average accuracy of 96% on a standard dataset. To investigate the underlying causes of the remaining 4% of recognition errors, we conducted a detailed analysis of misclassified samples across six standard datasets, as illustrated in Fig. 7 and Fig. 8. While previous research has typically categorized common benchmarks into regular and irregular text. However, these error samples indicates that the majority of incorrectly recognized text is not irregular. This suggests that, under the current training paradigm using large-scale real-world datasets, a more rigorous manual screening process for common benchmarks is warranted.
| Blurred | Artistic | Incomplete | Other | Total | Labelerr | |
| IIIT5k | 0 | 16 | 1 | 4 | 21 | 4 |
| SVT | 4 | 4 | 4 | 0 | 12 | 0 |
| ICDAR 2013 | 2 | 2 | 4 | 2 | 10 | 2 |
| ICDAR 2015 | 48 | 19 | 42 | 13 | 122 | 35 |
| SVTP | 7 | 6 | 12 | 7 | 32 | 4 |
| CUTE80 | 0 | 1 | 0 | 0 | 1 | 1 |
| Total | 61 | 48 | 63 | 26 | 198 | 46 |
| 30.81% | 24.24% | 31.82% | 13.13% | 100% |
Consequently, we identified five primary causes of recognition errors in these samples: (1) blurred, (2) artistic, (3) incomplete text, (4) image text labeling inconsistency, and (5) others. Specifically, the blurring text category encompasses issues such as low resolution, motion blur, or extreme lighting conditions. The artistic text category includes unconventional print fonts, predominantly found in business signage, as well as a proportion of handwritten text. Incomplete text refers to characters obscured by other objects or missing due to improper cropping, where missing information must be inferred from context. Image text labeling inconsistency means that there is an error in the given text label or there are some characters with phonetic symbols. As shown in Tab. 10, after excluding samples where the errors were due to labeling inconsistency, the remaining errors could be attributed to blurred (30.81%), artistic (24.24%), and incomplete text (31.82%), respectively. This classification allows us to conclude that SVTRv2’s recognition performance, particularly in complex scenarios involving blurred, artistic, or incomplete text, requires further enhancement.
12 Standardized Model Training Settings
The optimal hyperparameters for training different models often vary and are not universally fixed. However, critical factors such as training epochs, data augmentation techniques, input size, data type, and evaluation protocols have a substantial influence on model accuracy. To ensure fair and unbiased performance comparisons across models, these factors must be strictly standardized. Accordingly, we adopt a uniform training and evaluation setting, as shown in Table 11, to maintain consistency across all settings while simultaneously enabling each model to achieve its best possible accuracy.
| Setting | Detail |
| Training Set | For training, when the text length of a text image exceeds 25, samples with text length are randomly selected from the training set to ensure models are only exposed to short texts (length ). |
| Test Sets | For all test sets except the long-text test set (LTB), text images with text length are filtered. Text length is calculated by removing spaces and non-94-character-set special characters. |
| Input Size | Unless a method explicitly requires a dynamic size, models use a fixed input size of . If a model performs incorrectly with during training, the original size is used. The test input size matches the training size. |
| Data Augmentation | All models use the data augmentation strategy employed by PARSeq. |
| Training Epochs | Unless pre-training is required, all models are trained for 20 epochs. |
| Optimizer | AdamW is the default optimizer. If training fails to converge with AdamW, Adam or other optimizers are used. |
| Batch Size | Maximum batch size for all models is 1024. If single-GPU training is not feasible, 2 GPUs (512 per GPU) or 4 GPUs (256 per GPU) are used. If 4-GPU training runs out of memory, the batch size is halved, and the learning rate is adjusted accordingly. |
| Learning Rate | Default learning rate for batch size 1024 is 0.00065. The learning rate is adjusted multiple times to achieve the best results. |
| Learning Rate Scheduler | A linear warm-up for 1.5 epochs is followed by a OneCycle scheduler. |
| Weight Decay | Default weight decay is 0.05. NormLayer and Bias parameters have a weight decay of 0. |
| Data Type | All models are trained with mixed precision. |
| EMA or Similar Tricks | No EMA or similar tricks are used for any model. |
| Evaluation Protocols | Word accuracy is evaluated after filtering special characters and converting all text to lowercase. |